Find the matrix representation of T relative to the bases B and CFind the matrix representation of T relative to the bases B and C T: P2 +C, T(a + bx+ cx) = a+b+c a+b-c a-b+c B={1, x, x?}, C
Answer #1

a+b+c T: PB Ta+batex) = a+b-c a-btc T(1) T(4to.xto.x) 4 4 4 1 11 + O 0 7) T01 + 1. x+ 0.1) = 1 •3-8---) 10 TX) = Tod+On+dit